Find the sum of the first 6 terms of the Geometric progression : 18,6,2, .............
Answer by Boreal(15235)

common ratio is 1/3 by inspection
s=a1(1-r^n)/(1-r)
s=18*(1-1/729)/(2/3)
s=18(728/729)/(2/3)
18*3/2, invert divisor=27
27*(728/729), 27 divides into 729 27 times
it is 26 and 26/27
terms are 18,6,2,(2/3),(2/9),and (2/27)
They add to 26 26/27